OpenStudy (anonymous):

solve 4(a+9)=7[a-(3-a)]

OpenStudy (the_fizicx99):

Distribute! 4a + 36 = 7(2a-3) Can you solve from there? What I did was combine like terms in the second part of the equation. 7[a(3-a)

OpenStudy (anonymous):

is 40a=14-21 is this it i think

OpenStudy (the_fizicx99):

Where did you get 40a? ._.

OpenStudy (anonymous):

well i added them crap i did it wrong

OpenStudy (the_fizicx99):

You can't add like that lol xD only LIKE TERMS. Ex: 3a + 2a = 5a @dalia_lam please do not give out answers.

OpenStudy (anonymous):

ok thank you for helping me

OpenStudy (the_fizicx99):

Your next step is to multiply 7*2a and 7*(-3)

OpenStudy (the_fizicx99):

4a + 36 = 14a - 21 Can you solve now?

OpenStudy (anonymous):

40a=-7 whoo hoo i did it

OpenStudy (the_fizicx99):

No. @dalia_lam your answer is incorrect.

OpenStudy (the_fizicx99):

4a + 36 = 14a - 21 +21 +21 4a +57 = 14a -4a -4a 57 = 10a /10 /10 Whats the answer?

OpenStudy (the_fizicx99):

Did you understand?

OpenStudy (anonymous):

sorry, i thought it was -x

OpenStudy (the_fizicx99):

x?

OpenStudy (anonymous):